2.You know that the weight of an object is different on different planets. The weight of an object also varies based on the elevation on the elevation. The weight of an object is modeled: w=Cr^-2
Where C is a constant, and r is the distance that the object is from the center of the earth.
A. I weigh 120 pounds at sea level. If sea level is 3,950 miles from the center of the earth, determine the value of C that makes the equation true.
MY ANSWER, C=1872300000
Use the value of C you found in the previous question to determine how much I would weigh in
i. Death Valley (282 feet below Sea level)
ii. Alamagordo (4334 feet above sea level)

Answer by Boreal(15235)

w=C/r^2
C=w*r^2
=120*3950^2
C=1.8723 x 10^9
At Death Valley
282/5280 = =0.0534 miles
That is 3949.9466 miles
w=C/r^2=120.003 pounds
=================
Alamogordo is 3950.8208 miles from the center of the Earth.
w=1872300000/(3950.8208)^2=119.97 pounds